# Associate Everything Detected: Facilitating Tracking-by-Detection to the Unknown

### v1

@@ Abstract
Multi-object tracking (MOT) emerges as a pivotal and highly promising branch in the field of computer vision. Classical closed-vocabulary MOT (CV-MOT) methods aim to track objects of predefined categories. Recently, some open-vocabulary MOT (OV-MOT) methods have successfully addressed the problem of tracking unknown categories. However, we found that the CV-MOT and OV-MOT methods each struggle to excel in the tasks of the other. In this paper, we present a unified framework, Associate Everything Detected (AED), that simultaneously tackles CV-MOT and OV-MOT by integrating with any off-the-shelf detector and supports unknown categories. Different from existing tracking-by-detection MOT methods, AED gets rid of prior knowledge (e.g. motion cues) and relies solely on highly robust feature learning to handle complex trajectories in OV-MOT tasks while keeping excellent performance in CV-MOT tasks. Specifically, we model the association task as a similarity decoding problem and propose a sim-decoder with an association-centric learning mechanism. The sim-decoder calculates similarities in three aspects: spatial, temporal, and cross-clip. Subsequently, association-centric learning leverages these threefold similarities to ensure that the extracted features are appropriate for continuous tracking and robust enough to generalize to unknown categories. Compared with existing powerful OV-MOT and CV-MOT methods, AED achieves superior performance on TAO, SportsMOT, and DanceTrack without any prior knowledge. Our code is available at https://github.com/balabooooo/AED.
@@ I Introduction
Closed-vocabulary multi-object tracking (CV-MOT) methods [1, 2, 3, 4, 5] are designed to track a predefined, limited set of target classes consistent with training data. They have achieved remarkable results in CV-MOT datasets [6, 7, 8, 9]. Nevertheless, some practical applications like autonomous driving and augmented reality (AR) require trackers to handle new or unexpected classes. CV-MOT methods cannot expand the categories during the detection phase. In contrast, with the help of zero-shot capability of Multimodal Large Language Model (MLLM) [10, 11, 12], open-vocabulary multi-object tracking (OV-MOT) methods [13, 14, 15] can adapt to a broader range of categories, including unseen ones in training. However, when tracking certain categories, OV-MOT methods are less effective than well-finetuned CV-MOT methods.
@@
2) Some appearance-based methods [2, 4, 20] leverage the learned re-identification (ReID) features to compensate for the drawback of appearance-independent methods in CV-MOT (Fig. 1(b)). In particular, besides using prior knowledge, they train their appearance models or branches following the ReID learning pipeline (e.g. classification or triplet supervision [21]). Other methods [22, 13, 14] (Fig. 1(c)), especially for OV-MOT methods, use dual contrastive learning strategy for feature learning between two frames. However, the learning objectives of both ReID-based and dual contrastive-based methods are inconsistent with the inference objectives of the association stage for MOT. The association in MOT requires temporal discriminability, while ReID-based methods enable the model to have highly global identity (ID) discriminability by unsuitably modeling the association problem as a classification or alignment task within feature space. Dual contrastive-based methods lack temporal diversity, i.e., the association stage of MOT usually needs to handle intricate associations across far more than two frames. We argue that ignorance or lack of association-driven learning may lead to ambiguity during training.
@@
Figure 1: Motivation of AED. Existing closed-vocabulary and open-vocabulary trackers focus on tracking certain categories (e.g. person) and arbitrary categories (e.g. person, elephant, chair, bucket, etc) respectively. For the association phase, (a) appearance-independent methods rely on prior knowledge; (b) some appearance-based CV-MOT methods train their appearance branch following the ReID pipeline; (c) some other appearance-based methods use dual contrastive strategy for feature learning; (d) our proposed method unifies CV-MOT and OV-MOT tasks and uses association-centric learning.
@@
Figure 2: Overview of AED. During inference, a replaceable detector outputs boxes of arbitrary categories (e.g. cat and dog), and boxes are encoded into object queries. A CNN and a transformer encoder are responsible for extracting features of the frame. The sim-decoder calculates the similarities between object queries and track queries. Similarities are then used for matching. If a track query is matched (dashed lines), it will be updated by its object query. Matching results will be the track queries for the next frame.
@@ II-A Tracking-by-Detection MOT
In the past few years, tracking-by-detection has been a popular and high-performance paradigm for CV-MOT tasks. Objects are first detected and then associated. Guided by this, many methods have introduced innovative approaches in various aspects, such as motion features [1, 16, 17, 19, 23], appearance features [2, 20], and the real-time performance [3, 4] of the algorithms. SORT [1] introduces a Kalman filter [18] to predict the most probable location of objects in the next frame, which effectively solves the association problem. Based on SORT, DeepSORT [2] uses an extra re-identification (ReID) model to compensate for the shortcomings of IoU (Intersection over Union) distance. JDE [3] modifies Yolov3 [24] to simultaneously output detection results and ReID features. CSTrack [4] finds the competition between detection and ReID and further proposes a reciprocal network to learn task-dependent representations. ByteTrack [16] uses a simple but effective strategy to recover the true positive boxes with low confidence scores. However, these tracking-by-detection algorithms heavily rely on prior knowledge, especially motion cues (e.g. IoU distance). Some other advanced methods [17, 19, 20, 23] design handcrafted tracking logic (e.g. observation re-update and iterative scale-up IoU) to adapt to specific tracking scenarios such as dancing [9], pedestrian walking [7, 8], and sports [25]. These handcrafted tracking logics are not conducive to open-vocabulary tracking because motion patterns are hard to describe in scenarios with numerous categories. The reliance on handcrafted tracking logic significantly harms their universality in diverse scenarios.
@@ II-C Open-Vocabulary MOT
With the advancement of the MOT community, various MOT datasets [7, 8, 9] have abundant scenarios and a variety of motion patterns. However, these datasets predominantly feature the person category, which is not in accord with the practical and universal scenario. To address this problem, OV-MOT aims to track unseen categories. Unfortunately, very few papers have tackled the OV-MOT problem. Li et al. [13] first define the task of OV-MOT and propose OVTrack to handle this problem on the TAO [34] dataset. Different from classical CV-MOT (i.e. the categories for testing and training are consistent), OV-MOT splits categories of objects into non-overlapped $\mathcal{C}^{\mathrm{base}}$ and $\mathcal{C}^{\mathrm{novel}}$. Only $\mathcal{C}^{\mathrm{base}}$ is used for training and $\mathcal{C}^{\mathrm{base}}$ together with $\mathcal{C}^{\mathrm{novel}}$ are then used for testing by given names of the novel categories. OVTrack leverages distilled knowledge from CLIP [10] to track unseen categories $\mathcal{C}^{\mathrm{novel}}$. Moreover, it employs a contrastive learning approach to distinguish different IDs. By integrating 10 million automatically labeled data, GLEE [14] uses a unified framework to address several open-vocabulary tasks like detection, segmentation, grounding, tracking, etc. However, these methods overlook the true needs of the model during the association process leading to their suboptimal performance. NetTrack [15] leverages fine-grained features, i.e., point-level visual cues, to track highly dynamic objects. Nevertheless, its association performance is still less than ideal in some extremely dynamic scenarios like TAO.
@@ III Method
Figure 3: Design of (a) similarity decoder (sim-decoder) and (b) multi-head weight attention. Object queries and feature maps of the frame are fed into multi-head deformable cross-attention to extract features of objects. Then, the feed-forward network updates the object queries. Residual connection is omitted for simplicity. We use a multi-head weight cross-attention to calculate temporal similarities $S_{\mathrm{temporal}}$, which is then used for matching. During training, we use three multi-head weight attentions with shared weights to get temporal similarities $S_{\mathrm{temporal}}$ and weights $W_{\mathrm{temporal}}$, spatial similarities $S_{\mathrm{sptial}}$ and weights $W_{\mathrm{sptial}}$, and cross-clip similarities $S_{\mathrm{cross-clip}}$ and weights $W_{\mathrm{cross-clip}}$ respectively. These three similarity and weight matrices are then used for the association-centric learning mechanism.
@@ III-A Overview
We design AED to decode similarities between detection results and historical trajectories and perform matching. The input of AED is a video sequence. For every frame, there is an object detector (e.g., Co-DETR [35], YOLOX [36] or RegionCLIP [11]) outputting bounding boxes and a CNN backbone (i.e., ResNet50 [37]) followed with an encoder [32] extracting image features. For every frame, we encode the detected boxes into object queries. Therefore, we regard the detections as object queries and the existing trajectories as track queries of the current frame. Then the object queries and track queries are fed into a novel sim-decoder (Section III-B) to decode their similarities. To minimize the reliance on prior knowledge and adapt to the vast majority of motion patterns, AED leverages only the decoded similarity to accomplish the matching process.
Different from some existing tracking-by-query methods [31, 33] that track one trajectory by using one query, in AED, the matched track queries will be updated by the corresponding object queries (dashed lines in Fig. 2). Then, the matched object queries and the unmatched track queries are treated as track queries for the next frame. Note that we do not perform matching at the end of the first frame because no track query exists at that time.
To maximize the consistency between training and association and also enhance the robustness of AED to generalize to unknown categories, we propose an association-centric learning mechanism (Section III-C) during training. By leveraging the sim-decoder, the association-centric learning mechanism is reflected in three aspects: spatial, temporal, and cross-clip. Specifically, we apply spatial contrastive learning to make AED better distinguish objects within one single frame, which is the precondition for cross-frame associations. Then we take a sampled video clip with $n$ frames as the input and use temporal contrastive learning to help AED match objects with historical trajectories. This prompts AED to learn complex matching scenarios similar to the association phase. Moreover, we use cross-clip contrastive learning to improve the long-term ID consistency of AED on several frames. AED is only trained on specific categories following [13], while AED aims to associate arbitrary detections during inference, which is different from [13].
@@ III-B Similarity Decoder
As illustrated in Fig. 3(a), we propose a similarity decoder (sim-decoder) with $N$ layers to decode spatial, temporal, and cross-clip similarities $S$ and weight responses $W$. During inference, only $S_{\mathrm{temporal}}$ is required to perform matching. During training, all of the $S$ and $W$ are the input of association-centric learning which will be discussed in Section III-C.
@@
The input of the sim-decoder consists of three parts: object queries, track queries, and features map of the current frame. With the confidence scores available, we first encode $M_{t}$ bounding boxes from frame $t$ into object queries of size $M_{t}\times d_{\text{model}}$ by using sine-cosine positional encoding [26]. Similar to the inference phase, we also perform matching [38] during training to output real track queries for the next frame, while the input data is a sampled clip consisting of $n$ frames. At the beginning of the sim-decoder, a multi-head deformable cross-attention [32] is employed to aggregate the appearance features of every object query on the feature map. Then we apply a feed-forward network to update the object queries. To help sim-decoder better locate the feature, we employ box refinement using updated object queries following [33] (omitted in Fig. 3(a) for simplicity and will be discussed in Section III-D).
Then, we utilize the updated object queries from three dimensions during training. Firstly, as shown in Fig. 3(a), they are the queries $Q_{1}$ for the proposed multi-head weight cross-attention, and the track queries are the keys $K_{1}$. The output of multi-head weight cross-attention constitutes $S_{\mathrm{temporal}}$ and $W_{\mathrm{temporal}}$. They are the similarities and weight responses between current object queries and track queries. Their number of rows equals the number of object queries and the number of columns equals the number of track queries. $S_{\mathrm{temporal}}$ is then used for matching [38]. Secondly, they are the keys $K_{2}$ and queries $Q_{2}$ of multi-head weight self-attention to output $S_{\mathrm{spatial}}$ and $W_{\mathrm{spatial}}$, which are similarities and weight responses among object queries from the present frame. The number of rows and columns of $S_{\mathrm{spatial}}$ and $W_{\mathrm{spatial}}$ both equals the number of object queries in the current frame. Thirdly, we store all of the output object queries from the sampled clip in a clip query buffer during training. At the end of the last frame, we take all object queries from the buffer as both keys $K_{3}$ and queries $Q_{3}$ for multi-head weight self-attention to output $S_{\mathrm{cross-clip}}$ and $W_{\mathrm{cross-clip}}$. They denote the similarity and weight responses among all the objects in the clip. Their number of rows and columns also equals the total number of objects. Note that the three mentioned multi-head weight attention share the same weights. All of the $S$ and $W$ are further used for association-centric learning and will be discussed in Sec. III-C.
@@
We further introduce our proposed multi-head weight attention. Inspired by the ”weight” concept in multi-head attention [26], our multi-head weight attention is only responsible for calculating the similarities $S$ and weight responses $W$ between the input keys $K$ and queries $Q$. As illustrated in Fig. 3(b), we simplify the multi-head attention to specifically compute the $S$ and $W$ between queries and keys and abandon the concept of ”value” in multi-head attention. The more similar a given pair of the key and query (e.g. from the same ID), the higher the corresponding similarity and response will be. If $K=Q$, it is called multi-head weight self-attention, otherwise, it is called cross-attention. To be specific, we linearly project $K$ and $Q$ for $h$ times respectively, where $h$ is the number of heads. For each $\text{head}_{i}$, we compute the similarities $S_{i}$ and the weight responses $W_{i}$ between the projected queries and keys using weight attention. $h$ weight attentions are employed for the similarity learning from different representation subspaces. For $\text{head}_{i}$, its $S_{i}$ and $W_{i}$ can be formulated as:
@@
Different from [26], for $S_{i}$, we employ cosine similarity instead of dot product and softmax in our weight attention because the softmax function can only identify the relative maximum, not the absolute maximum, which is catastrophic for data association in MOT. For example, newly emerged objects should have low similarity with all historical trajectories. If softmax is used, the similarities will be comparatively high.
The results of $h$ weight attentions are then averaged element-wise to output the final $S$ and $W$ for multi-head weight attention:
@@
During inference, we take the entire video sequence as the input. Instead of leveraging some prior knowledge like Kalman filter [18], we directly use the $S_{\mathrm{temporal}}$ from the final layer of sim-decoder as the input for the Hungarian algorithm [38] to perform matching. Again, we simply update the track query with the corresponding object query if they match successfully. Inspired by [16], we employ a very simple two-stage matching strategy. Specifically, We first associate the high-scoring boxes and then proceed to associate the low-scoring boxes. AED inits every untracked bounding box as a new trajectory without any complex initialization procedures like [2, 16, 13].

### v2

@@ Abstract
Multi-object tracking (MOT) emerges as a pivotal and highly promising branch in the field of computer vision. Classical closed-vocabulary MOT (CV-MOT) methods aim to track objects of predefined categories. Recently, some open-vocabulary MOT (OV-MOT) methods have successfully addressed the problem of tracking unknown categories. However, we found that the CV-MOT and OV-MOT methods each struggle to excel in the tasks of the other. In this paper, we present a unified framework, Associate Everything Detected (AED), that simultaneously tackles CV-MOT and OV-MOT by integrating with any off-the-shelf detector and supports unknown categories. Different from existing tracking-by-detection MOT methods, AED gets rid of prior knowledge (e.g., motion cues) and relies solely on highly robust feature learning to handle complex trajectories in OV-MOT tasks while keeping excellent performance in CV-MOT tasks. Specifically, we model the association task as a similarity decoding problem and propose a sim-decoder with an association-centric learning mechanism. The sim-decoder calculates similarities in three aspects: spatial, temporal, and cross-clip. Subsequently, association-centric learning leverages these threefold similarities to ensure that the extracted features are appropriate for continuous tracking and robust enough to generalize to unknown categories. Compared with existing powerful OV-MOT and CV-MOT methods, AED achieves superior performance on TAO, SportsMOT, and DanceTrack without any prior knowledge. Our code is available at https://github.com/balabooooo/AED.
@@ I Introduction
Closed-vocabulary multi-object tracking (CV-MOT) methods [1, 2, 3, 4, 5] are designed to track a predefined, limited set of target classes consistent with training data. They have achieved remarkable results in CV-MOT datasets [6, 7, 8, 9]. Nevertheless, some practical applications like autonomous driving [10] and augmented reality (AR) require trackers to handle new or unexpected classes. CV-MOT methods cannot expand the categories during the detection phase. In contrast, with the help of zero-shot capability of Multimodal Large Language Model (MLLM) [11, 12, 13], open-vocabulary multi-object tracking (OV-MOT) methods [14, 15, 16, 17] can adapt to a broader range of categories, including unseen ones in training. However, when tracking certain categories, OV-MOT methods are less effective than well-finetuned CV-MOT methods.
@@
2) Some appearance-based methods [2, 4, 22] leverage the learned re-identification (ReID) features to compensate for the drawback of appearance-independent methods in CV-MOT (Fig. 1(b)). In particular, besides using prior knowledge, they train their appearance models or branches following the ReID learning pipeline (e.g., classification or triplet supervision [23]). Other methods [24, 14, 15] (Fig. 1(c)), especially for OV-MOT methods, use dual contrastive learning strategy for feature learning between two frames. However, the learning objectives of both ReID-based and dual contrastive-based methods are inconsistent with the inference objectives of the association stage for MOT. The association in MOT requires temporal discriminability, while ReID-based methods enable the model to have highly global identity (ID) discriminability by unsuitably modeling the association problem as a classification or alignment task within feature space. Dual contrastive-based methods lack temporal diversity, i.e., the association stage of MOT usually needs to handle intricate associations across far more than two frames. We argue that ignorance or lack of association-driven learning may lead to ambiguity during training.
@@
Figure 1: Motivation of AED. Existing closed-vocabulary and open-vocabulary trackers focus on tracking certain categories (e.g., person) and arbitrary categories (e.g., person, elephant, chair, bucket, etc) respectively. For the association phase, (a) appearance-independent methods rely on prior knowledge; (b) some appearance-based CV-MOT methods train their appearance branch following the ReID pipeline; (c) some other appearance-based methods use dual contrastive strategy for feature learning; (d) our proposed method unifies CV-MOT and OV-MOT tasks and uses association-centric learning.
@@
Figure 2: Overview of AED. During inference, a replaceable detector outputs boxes of arbitrary categories (e.g., cat and dog), and boxes are encoded into object queries. A CNN and a transformer encoder are responsible for extracting features of the frame. The sim-decoder calculates the similarities between object queries and track queries. Similarities are then used for matching. If a track query is matched (dashed lines), it will be updated by its object query. Matching results will be the track queries for the next frame. We use different shapes to represent different IDs and different colors within each shape represent frames from different times.
@@ II-A Tracking-by-Detection MOT
In the past few years, tracking-by-detection has been a popular and high-performance paradigm for CV-MOT tasks. Objects are first detected and then associated. Guided by this, many methods have introduced innovative approaches in various aspects, such as motion features [1, 18, 19, 21, 25], appearance features [2, 22], and the real-time performance [3, 4] of the algorithms. SORT [1] introduces a Kalman filter [20] to predict the most probable location of objects in the next frame, which effectively solves the association problem. Based on SORT, DeepSORT [2] uses an extra re-identification (ReID) model to compensate for the shortcomings of IoU (Intersection over Union) distance. JDE [3] modifies Yolov3 [26] to simultaneously output detection results and ReID features. CSTrack [4] finds the competition between detection and ReID and further proposes a reciprocal network to learn task-dependent representations. ByteTrack [18] uses a simple but effective strategy to recover the true positive boxes with low confidence scores. However, these tracking-by-detection algorithms heavily rely on prior knowledge, especially motion cues (e.g., IoU distance). Some other advanced methods [19, 21, 22, 25] design handcrafted tracking logic (e.g., observation re-update and iterative scale-up IoU) to adapt to specific tracking scenarios such as dancing [9], pedestrian walking [7, 8], and sports [27]. These handcrafted tracking logics are not conducive to open-vocabulary tracking because motion patterns are hard to describe in scenarios with numerous categories. The reliance on handcrafted tracking logic significantly harms their universality in diverse scenarios.
@@ II-C Open-Vocabulary MOT
With the advancement of the MOT community, various MOT datasets [7, 8, 9] have abundant scenarios and a variety of motion patterns. However, these datasets predominantly feature the person category, which is not in accord with the practical and universal scenario. To address this problem, OV-MOT aims to track unseen categories. Unfortunately, very few papers have tackled the OV-MOT problem. Li et al. [14] first define the task of OV-MOT and propose OVTrack to handle this problem on the TAO [36] dataset. Different from classical CV-MOT (i.e., the categories for testing and training are consistent), OV-MOT splits categories of objects into non-overlapped $\mathcal{C}^{\mathrm{base}}$ and $\mathcal{C}^{\mathrm{novel}}$. Only $\mathcal{C}^{\mathrm{base}}$ is used for training, and $\mathcal{C}^{\mathrm{base}}$ together with $\mathcal{C}^{\mathrm{novel}}$ is then used for testing by providing names of the novel categories. OVTrack leverages distilled knowledge from CLIP [11] to track unseen categories $\mathcal{C}^{\mathrm{novel}}$. Moreover, it employs a contrastive learning approach to distinguish different IDs. By integrating 10 million automatically labeled data, GLEE [15] uses a unified framework to address several open-vocabulary tasks like detection, segmentation, grounding, tracking, etc. However, these methods overlook the true needs of the model during the association process leading to their suboptimal performance. NetTrack [16] leverages fine-grained features, i.e., point-level visual cues, to track highly dynamic objects. Nevertheless, its association performance is still less than ideal in some extremely dynamic scenarios like TAO.
@@ III Method
Figure 3: Design of (a) similarity decoder (sim-decoder) and (b) multi-head weight attention. Object queries and feature maps of the current frame are fed into a multi-head deformable cross-attention and feed-forward network to extract appearance features of object queries, where the residual connection is omitted for simplicity. During training, we use three multi-head weight attentions (i.e., Temporal/Spatial/Cross-Clip Attention) with shared weights to get temporal similarities $S_{\mathrm{temporal}}$ and weights $W_{\mathrm{temporal}}$, spatial similarities $S_{\mathrm{sptial}}$ and weights $W_{\mathrm{sptial}}$, and cross-clip similarities $S_{\mathrm{cross-clip}}$ and weights $W_{\mathrm{cross-clip}}$ respectively. The temporal similarities $S_{\mathrm{temporal}}$ is then used for matching. These three similarity and weight matrices are then used for the association-centric learning mechanism.
@@ III-A Overview
We design AED to decode similarities between detection results and historical trajectories and perform matching. The input of AED is a video sequence. For every frame, there is an object detector (e.g., Co-DETR [37], YOLOX [38] or RegionCLIP [12]) outputting bounding boxes and a CNN backbone (i.e., ResNet50 [39]) followed with an encoder [34] extracting image features. For every frame, we encode the detected boxes into object queries. Therefore, we regard the detections as object queries and the existing trajectories as track queries of the current frame. Then the object queries and track queries are fed into a novel sim-decoder (Section III-B) to decode their similarities. The sim-decoder also aggregates the appearance features of object queries layer by layer. To minimize the reliance on prior knowledge and adapt to the vast majority of motion patterns, AED leverages only the decoded similarity to accomplish the matching process.
Different from existing tracking-by-query methods [33, 35] that track one trajectory by using one query, in AED, the matched track queries will be updated by the corresponding object queries. Then, the matched object queries and the unmatched track queries are treated as track queries for the next frame. Note that we do not perform matching at the end of the first frame because no track query exists at that time.
To maximize the consistency between training and online association and also enhance the robustness of AED to generalize to unknown categories and motion patterns, we propose an association-centric learning mechanism (Section III-C) during training. Association-centric learning prompts AED to learn complex matching scenarios that are similar to the association phase. By leveraging the sim-decoder, the association-centric learning mechanism is reflected in three aspects: spatial, temporal, and cross-clip. Specifically, we apply spatial contrastive learning to make AED better distinguish objects within one single frame, which is the precondition for cross-frame associations. Then we take a sampled video clip with $n$ frames as the input and use temporal contrastive learning to help AED match objects with historical trajectories. Moreover, we use cross-clip contrastive learning to improve the long-term ID consistency of AED on several frames.
@@ III-B Similarity Decoder
We formulate the learning process of AED as a similarity and weight response decoding problem. As illustrated in Fig. 3(a), we propose a similarity decoder (sim-decoder) with $N$ layers to decode spatial, temporal, and cross-clip similarities $S$ and weight responses $W$.
@@
The input of the sim-decoder consists of three parts: object queries, track queries, and features map of the current frame. Based on the confidence scores output by the detector, we first encode $M_{t}$ bounding boxes from frame $t$ into object queries of size $M_{t}\times d_{\text{model}}$ by using sine-cosine positional encoding [28], where $d_{\text{model}}$ denotes the feature dimension of each query. Since the object queries only contain confidence information at this time, we employ a multi-head deformable cross-attention [34] and a feed-forward network to aggregate their appearance features according to the feature map. To help sim-decoder better locate the feature, we employ box refinement using object queries following [35] (omitted in Fig. 3(a) for simplicity and will be discussed in Section III-D). We further utilize object queries in three attentional dimensions during training.
III-B1 Temporal Attention
Temporal attention takes track queries and object queries as the input keys ($K_{1}$) and queries ($Q_{1}$). The track queries consist of the matched object queries and the unmatched track queries from the previous frame. This attention calculates pairwise similarities $S_{\mathrm{temporal}}$ and weight responses $W_{\mathrm{temporal}}$ between object and track queries. $S_{\mathrm{temporal}}$ and $W_{\mathrm{temporal}}$ are then employed in association-centric learning (Section III-C) to guarantee robust matching between detections and trajectories.
III-B2 Spatial Attention
The spatial attention then operates on the object queries within a frame to compute pairwise $S_{\mathrm{spatial}}$ and $W_{\mathrm{spatial}}$ between these queries. These matrices physically represent the similarity measurements and weights among detections within individual frames. $S_{\mathrm{spatial}}$ and $W_{\mathrm{spatial}}$ then help to keep the spatial discriminability of AED in association-centric learning.
III-B3 Cross-Clip Attention
During training, all output object queries from the sampled video clip are cached in a clip query buffer. Upon reaching the terminal frame of the clip, these buffered queries are simultaneously assigned as keys ($K_{3}$) and queries ($Q_{3}$) for cross-clip attention computation, generating $S_{\mathrm{cross-clip}}$ and $W_{\mathrm{cross-clip}}$. These matrices denote the similarities and weight responses among all the objects that appear in this clip. Through association-centric learning, $S_{\mathrm{cross-clip}}$ and $W_{\mathrm{cross-clip}}$ are further optimized to improve the long-term trajectory modeling capability of AED.
@@ III-B3 Cross-Clip Attention
All three attention mechanisms mentioned above are based on our proposed multi-head weight attention, as illustrated in Fig. 3(b). Inspired by the ”weight” concept in standard multi-head attention [28], we reformulate it as multi-head weight attention by eliminating the ”value” component, specializing in deriving similarities $S$ and weight responses $W$ between input keys $K$ and queries $Q$. The more similar a given pair of the key and query (e.g., from the same ID), the higher the corresponding similarity and response will be. To be specific, we linearly project $K$ and $Q$ for $h$ times respectively, where $h$ is the number of heads. For each $\text{head}_{i}$, we compute the similarities $S_{i}$ and the weight responses $W_{i}$ between the projected queries and keys using weight attention. $h$ weight attentions are employed for the similarity learning from different representation subspaces. For $\text{head}_{i}$, its $S_{i}$ and $W_{i}$ can be formulated as:
@@
The elements in $S_{i}$ and $W_{i}$ are the cosine similarities and dot products between linearly mapped keys and queries respectively. Different from [28], for $S_{i}$, we employ cosine similarity instead of dot product with softmax in our weight attention because the softmax function can only identify the relative maximum, not the absolute maximum, which is catastrophic for data association in MOT. For example, newly emerged objects should have low similarity with all historical trajectories. If softmax is used, the similarities will be comparatively high.
The results of $h$ weight attentions are then averaged element-wise to output the final $S$ and $W$ for multi-head weight attention (i.e., temporal/spatial/cross-clip attention in Fig. 3(a)):
@@
During inference, only $S_{\mathrm{temporal}}$ is required to perform matching, and we take the entire video sequence as the input. Instead of leveraging some prior knowledge like Kalman filter [20], we directly use the $S_{\mathrm{temporal}}$ from the final layer of sim-decoder as the input for the Hungarian algorithm [40] to perform matching. Again, we simply update the track query with the corresponding object query if they match successfully. Inspired by [18], we employ a very simple two-stage matching strategy. Specifically, we first associate the high-scoring boxes and then proceed to associate the low-scoring boxes. AED inits every untracked bounding box as a new trajectory without any complex initialization procedures like [2, 18, 14].